Program Information
Recent Log Entries
- COMPLETED Troubleshoot North Access Cooling Tower Pump Controls
- COMMENT ON C75-05 (2L03) SRF Cryomodule Commissioning
- DecaRads are on 1L05, 1L10, 1L25
- COMMENT ON SL25 and SL26 vacuum controls and cable installation for reworked 2L25/2L26 girders
- COMMENT ON Post SAM FE onset measurements C75 and C100 cryomodules
- Gun HV Turned Off
- Gun HV Rampdown Initiated
- COMMENT ON Removal and Installation of Girders 2L25A & 2L26A
- COMMENT ON Plasma processing during the Fall 2025 SAM
- COMMENT ON Flash new firmware on all C-100 and GP Series Vacuum Control Chassis (SAM 2025-26)
- HYST AREA 7-0 Cycles Magnets
- HYST AREA 7-0 Cycles Magnets
- HYST AREA 7-0 Cycles Magnets
- COMPLETED Remove West Arc to BSY crossover beam pipe
- The following cavity histories have been updated: Zone 2L23











